# Split Unlearningthanks: Accepted by ACM CCS’25

### v6

@@ 1 Introduction
Split Learning (SL) is one of the latest approaches to distributed learning, suitable for scenarios where clients with limited computational resources, such as smartphones or IoT devices, need to offload part of their training to the servers [1, 2, 3]. Existing SL frameworks include two configurations: Vanilla SL and U-shaped SL [4]. In Vanilla SL, the model is split into two segments: The clients manage the segment receiving the training data and send intermediate values to the server, while the server handles the segment producing the final outputs without accessing raw data. U-shaped SL extends this by eliminating the need for label sharing, protecting both label privacy and data privacy. The model in U-shaped SL is divided into three segments: The clients process the input data and produce output classes, while the server manages the intermediate layers.
@@
The need for the compliance with regulatory requirements for “the right to forget” (e.g., GDPR [5]), which grants individuals the right to remove their personal data or impact of the data be erased, necessitates the development of effective machine unlearning [6] techniques. However, the requirement becomes challenging within the context of SL, where the model segments across clients and servers need to be jointly updated. Current methods include retraining the model [7, 8] or using alternatives like fine-tuning, gradient ascent, and knowledge distillation [9, 10, 11]. While these alternatives are favored for their higher unlearning efficiency (i.e. lower time consumption) compared to retraining, which prioritizes effectiveness (i.e., accuracy[12], they all require all data samples during each update iteration. This becomes problematic when some data samples are absent due to contextual constraints. Sharded, Isolated, Sliced, and Aggregated (SISA) methods [6, 13, 14] address this by segmenting data and training multiple models in parallel, maintaining higher effectiveness and efficiency than non-SISA unlearning methods.
Typical SL solutions [1, 2, 3, 15] are inadequate for directly applying SISA unlearning in multi-client scenarios, where multiple clients have independent models while sharing a common model on the server. Although clients’ training data can be considered “Sharded”, SL undermines the “Isolated” requirement of SISA because: (1) the server model is tightly coupled with the client models through frequent bidirectional propagation, and (2) the shared server model is iteratively trained across all clients. Thus, a new framework is needed to decouple the shared server model from the client models, enabling SISA unlearning in multi-client SL environments.
@@
[…]
@@ 2 Related Work
Unlearning in distributed environments. Machine unlearning has not been fully explored within current distributed learning frameworks, including SL. While efficient unlearning strategies have gained attention in Federated Learning (FL), which also involves collaboration between clients and a central server. Some researchers have worked on accelerating retraining across all clients, the most robust unlearning method in FL, using optimized algorithms like distributed Newton-type updates [16], SGA-EWC [17], and SFU [18]. Solutions like FedEraser [19] reduce the impact on non-unlearning-involved clients through calibration but still require cooperation from some clients. However, in SL, where client resources are often limited, non-requesting clients may lack the incentive or ability to contribute, especially if they are uninterested in or unable to participate in the unlearning process due to network issues, such as low bandwidth, high latency, or intermittent connectivity. This presents a significant challenge in SL compared to other distributed environments like FL, where developing unlearning methods that accommodate these constraints remains an unresolved issue.
@@
SISA guideline for SL unlearning. SISA [6] is a state-of-the-art guideline for robust and efficient machine unlearning, enabling rapid adaptation to data removal requests without requiring full model retraining. By strategically partitioning data into distinct “shards” linked to different model versions, the SISA guideline allows for the targeted unlearning of specific data shards without requiring the participation of all shards. This method has been successfully applied in various domains, including unlearning in graph neural networks [13] and large language models [20]. However, SISA unlearning faces significant challenges in multi-client SL scenarios due to the interdependence between client model segments and the shared server model segment, violating the Isolated requirement of SISA. Consequently, the direct application of conventional SISA is impractical for existing SL scenarios.
@@
Label privacy under SL unlearning. Beyond unlearning, privacy concerns regarding intermediate outputs and labels in SL have been widely studied. U-shaped SL [4] avoids label propagation to the server by placing the output layer on the client side but violates SISA’s Isolated requirement by necessitating the return of intermediate values across all clients, disabling the independent unlearning and adding overhead to resource-limited clients. If labels are shared but privacy is required, existing research primarily addresses privacy risks by directly adding noise and obfuscating labels [21, 22] or further protecting intermediate outputs transmitted to the server. Wu et al. [23] introduced a DP mechanism to perturb these transmitted outputs, while Li et al. [24] developed the Marvell method to quantify privacy leakage and optimize noise structures. However, these methods overlook the exposure of sensitive information through label quantity and semantics. Xiao et al. [25] obfuscated both labels and outputs via multiple linear combinations to reduce association with original data but still failed to conceal label quantity, potentially exposing sensitive information. Protecting the quantity of labels is critical in label-sharing SL, as revealing the exact number of labels allows adversaries to exploit supervised learning techniques that achieve higher accuracy by relying on this information, significantly increasing the risk of inference attacks and unauthorized model reconstruction.
@@
RRWithPrior [21]
LPSC [22]
DP-based Defense [23]
Marvell [24]
MALM [25]
U-shaped SL [4]
@@ 3.1 Split Learning
SL is a distributed ML technique designed to train neural networks without sharing raw data between $K$ clients, each owning datasets $D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$, and the server that holds no data [4, 26, 27]. A typical SL solution involves two key factors:
@@
The above processes enable training without sharing raw data or model details between clients and the server. In multi-client, single-server scenarios, two primary training methods are used: synchronizing model weights between clients after each epoch or alternating training epochs between clients and the server without synchronization [4].
Incremental training process. This process has been utilized by many learning frameworks, including FL [28, 29] and SL, where a model is continuously updated by gradually incorporating new data over time. Each iteration of back-propagation during model training needs to be sent back to every client involved in the training process and builds incrementally on the previous ones, embodying the incremental nature of SL [30]. This process enables the model to dynamically learn and adapt as new data becomes available.
The challenges of applying SISA in multi-client SL scenarios stem from the incremental training process inherent in these setups. This iterative process undermines the isolated requirement of SISA due to the tight coupling between the server and client model segments through regular bidirectional propagation and the shared server model being incrementally trained across all clients. The knowledge to be unlearned, reflected in $W_{k}^{(t)}$, would have also influenced $W_{k^{\prime}}^{(t^{\prime})}$ for all $k^{\prime}\neq k$ and $t^{\prime}\geq t$, necessitating the recalculation of corresponding changes in subsequent updates, thereby making it impossible to apply SISA unlearning [31, 12].
@@ 3.2 Machine Unlearning
Machine unlearning [32, 33, 10, 9] involves removing the influence of specific data subsets from a trained model while preserving its performance on the remaining data. The goal is to eliminate the impact of certain labels on the model’s predictions without compromising accuracy on the retained data. Let $D=\{\left(X_{i},Y_{q}\right)\}_{i=1,q=1}$ represent the dataset, where $X_{i}$ is the $i$-th training sample and $Y_{q}\in\{1,\cdots,Q\}$ is its label. Unlearning removes the influence of subset $D_{u}\subset D$ with labels $Y_{u}$ while preserving performance on the remaining data $D_{r}=D\backslash D_{u}$ with labels $Y_{o}$, ensuring $Y_{u}$ is forgotten and predictions $Y_{o}$ remain accurate.
Retraining. The most direct and effective, though not efficient, method of implementing unlearning is retraining [7, 8]. In this method, the revoked sample is deleted, and the model is retrained on the original dataset minus the deleted sample, denoted as $D_{r}=D\setminus{((X_{i},Y_{q})}\subset D_{u})$. While this method is effective and straightforward, the computational overhead becomes prohibitive for complex models and large datasets.
Alternatives. Referring to fine-tuning, gradient ascent, and knowledge distillation, These unlearning methods are proposed to improve unlearning efficiency compared to retraining, but they still consider retraining as the baseline for unlearning effectiveness [9, 10, 11], including:
@@
SISA. SISA [6] is a guideline designed to enable efficient and independent unlearning for those who request it, without affecting other participants. In SISA, the original training set $D_{o}$ is divided into $K$ disjoint shards {$D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$}, each used to train a separate model $\mathcal{F}_{o}^{1},\mathcal{F}_{o}^{2},\ldots,\mathcal{F}_{o}^{K}$}. These models generate individual predictions, which are then aggregated to produce a global prediction. When an unlearning request is made, only the shard containing the specific data sample needs to be retrained, allowing the unlearning process to be conducted independently for the requesting party, without impacting other clients or the overall system.
@@ 4.1 Problem Definition
    G1: Independent unlearning. This goal ensures that unlearning affects only the requesting client and server, leaving other clients ($k\textquoteright\neq k$) and their resources unaffected.
@@ 5 SplitWiper+: Privacy-Preserving Variation via Label Expansion
    Internal users. This includes clients ($\mathbb{C}_{inter}$) and servers ($\mathbb{S}_{inter}$) involved in training, modeled as honest-but-curious [24]. They adhere to the protocol without tampering, such as sending incorrect features or causing failures. Each client ${C}_{inter}\in\mathbb{C}_{inter}$ knows the shared labels among participants but keeps their complete label set $\mathbb{Y}_{inter}$ private. Clients also aim to hide the quantity and semantics of their labels from the server $\mathbb{S}_{inter}$. This model emphasizes privacy preservation in internal label sharing.
@@
    External attackers. An external attacker $\mathbb{A}_{ext}$ is considered to know the distribution of the original dataset fed into the input layer at the targeted client $\mathbb{C}_{target}$, and the posterior distribution of the model at the server side. The attacker $\mathbb{A}_{ext}$ aims to use a shadow dataset that mimics these distributions to construct an attack model for conducting inference attacks [34]. This player model focuses on the level of privacy preservation in the unlearning capability of the framework.
@@
Protecting label quantity matters. In G3, protecting the label privacy goes beyond protecting their semantics to include the quantity of labels, which is crucial for data privacy. By concealing the exact number of labels, adversaries are prevented from applying supervised learning methods that rely on knowing the number of classes for high accuracy. Instead, they are forced to use unsupervised learning techniques, lacking clear objectives and explicit class labels and generally resulting in lower accuracy [35, 36]. Protecting label quantity thus reduces the risk of unauthorized model reconstruction and inference attacks, strengthening the overall privacy and security of the learning process.
@@ 5.1 Label Expansion in SplitWiper+
Label expansion and anonymization (Line 13-16). Once the complete set $\mathbb{Y}$ is aggregated, clients collaborate to assign an expansion factor $\gamma_{q}$ to each label $Y_{q}$. This factor, $\gamma_{q}$, can be a randomly selected integer within a predefined range $\Gamma_{p}$ for each $Y_{q}$, or be a constant integer $\gamma>1$ and uniformly applied across all labels, i.e., $\gamma_{q}=\gamma,\forall q$. Each label, $Y_{q}$, is then expanded to $\gamma_{q}$ instances, forming the initial expanded label set $\mathbb{Y}_{exp}=\{Y_{1}^{*},\cdots,Y_{e}^{*},\cdots Y_{E}^{*}\}$, where $E=\sum_{q=1}^{Q}{\gamma_{q}}$. This expansion process masks the total count $Q$ of real labels, preventing the server from inferring specific client tasks or data types based on label quantity. The elements within $\mathbb{Y}_{exp}$ are shuffled and re-indexed to anonymize label semantics. While we use a basic pseudonymization technique, our framework can also accommodate methods like hashing [37] or encryption [38, 39] based on specific needs. The mapping between $\mathbb{Y}$ and $\mathbb{Y}_{exp}$ is preserved in $\mathcal{G}_{Y}$, ensuring that labels are expanded and anonymized without losing semantic integrity.

### v8

# Baihe Ma1, Caijun Sun3, Wei Ni1, Ren Ping Liu1

@@ 1 Introduction
Split Learning (SL) is one of the latest approaches to distributed learning, suitable for scenarios where clients with limited computational resources, such as smartphones or IoT devices, need to offload part of their training to the servers [poirot2019split, vepakomma2018split, thapa2022splitfed]. Existing SL frameworks include two configurations: Vanilla SL and U-shaped SL [GUPTA20181]. In Vanilla SL, the model is split into two segments: The clients manage the segment receiving the training data and send intermediate values to the server, while the server handles the segment producing the final outputs without accessing raw data. U-shaped SL extends this by eliminating the need for label sharing, protecting both label privacy and data privacy. The model in U-shaped SL is divided into three segments: The clients process the input data and produce output classes, while the server manages the intermediate layers.
@@
The need for the compliance with regulatory requirements for “the right to forget” (e.g., GDPR [voigt2017eu]), which grants individuals the right to remove their personal data or impact of the data be erased, necessitates the development of effective machine unlearning [sisa] techniques. However, the requirement becomes challenging within the context of SL, where the model segments across clients and servers need to be jointly updated. Current methods include retraining the model [retrain, retrain1] or using alternatives like fine-tuning, gradient ascent, and knowledge distillation [qu2023learn, nguyen2022survey, distil1Unl]. While these alternatives are favored for their higher unlearning efficiency (i.e. lower time consumption) compared to retraining, which prioritizes effectiveness (i.e., accuracy) [liu2024fishers], they all require all data samples during each update iteration. This becomes problematic when some data samples are absent due to contextual constraints. Sharded, Isolated, Sliced, and Aggregated (SISA) methods [sisa, chen2022graph, chen2022recommendation] address this by segmenting data and training multiple models in parallel, maintaining higher effectiveness and efficiency than non-SISA unlearning methods.
Typical SL solutions [poirot2019split, vepakomma2018split, thapa2022splitfed, zhang2023privacy] are inadequate for directly applying SISA unlearning in multi-client scenarios, where multiple clients have independent models while sharing a common model on the server. Although clients’ training data can be considered “Sharded”, SL undermines the “Isolated” requirement of SISA because: (1) the server model is tightly coupled with the client models through frequent bidirectional propagation, and (2) the shared server model is iteratively trained across all clients. Thus, a new framework is needed to decouple the shared server model from the client models, enabling SISA unlearning in multi-client SL environments.
@@
$\triangleright\quad$ We introduce SplitWiper (Sec.4), a new SL framework that implements the SISA guideline, achieving SISA unlearning capabilities: (i) selective involvement of only clients who request unlearning and (ii) effective unlearning while maintaining comparable utility for retained data. SplitWiper features an efficient and streamlined interaction between clients and the server by incorporating a new low-effort pre-training phase on clients, after which the clients’ weights are frozen (one-off). It allows iterative updates to focus only on the server’s weights without requiring feedback loops or updates to be returned to clients (one-way). This novel one-way-one-off protocol architecture for propagating neural signals from clients to the server forms the foundation for enabling SISA unlearning in SplitWiper. Our solution removes the requirement for all clients to engage in unlearning processes by Isolating training on each client, allowing for the formation of distinct Shards or Slices of data that are Aggregated exclusively on the server.
$\triangleright\quad$ We further introduce SplitWiper+ (Sec.5), a privacy-enhancing variation of SplitWiper, to address the privacy concerns associated with label sharing during the learning and unlearning tasks. The “Isolated” principle of SISA invalidates traditional privacy-preserving methods, such as non-label-sharing U-shaped SL [vepakomma2018split]. SplitWiper+ incorporates a novel label expansion strategy, marking the first approach to mask the quantity of training labels and their semantics. By expanding, shuffling, and anonymizing real labels, while protecting the outputs of clients’ models with differential privacy (DP), this strategy prevents adversaries from exploiting the exact number of labels and launching effective attacks with supervised learning techniques. This innovative design effectively mitigates privacy leakage when clients’ labels are exposed to the server during propagation.
$\triangleright\quad$ We conduct extensive experiments (Sec.6) in various settings where the clients either share the same training task with various data distributions or engage in different tasks, exploring the framework’s adaptability and performance across a range of scenarios. The results show that SplitWiper achieves complete unlearning accuracy (0%) while improving retained accuracy by 8%, engaging only the clients who request unlearning to minimize involvement and reduce overhead. SplitWiper maintains constant overhead, cutting off computational and communication costs by over 99% compared to existing SL frameworks. Additionally, SplitWiper+ preserves over 90% of client label privacy from the server.
@@ 2 Related Work
Unlearning in distributed environments. Machine unlearning has not been fully explored within current distributed learning frameworks, including SL. While efficient unlearning strategies have gained attention in Federated Learning (FL), which also involves collaboration between clients and a central server. Some researchers have worked on accelerating retraining across all clients, the most robust unlearning method in FL, using optimized algorithms like distributed Newton-type updates [9796721], SGA-EWC [9964015], and SFU [li2023subspace]. Solutions like FedEraser [liu2021federaser] reduce the impact on non-unlearning-involved clients through calibration but still require cooperation from some clients. However, in SL, where client resources are often limited, non-requesting clients may lack the incentive or ability to contribute, especially if they are uninterested in or unable to participate in the unlearning process due to network issues, such as low bandwidth, high latency, or intermittent connectivity. This presents a significant challenge in SL compared to other distributed environments like FL, where developing unlearning methods that accommodate these constraints remains an unresolved issue.
@@
SISA guideline for SL unlearning. SISA [sisa] is a state-of-the-art guideline for robust and efficient machine unlearning, enabling rapid adaptation to data removal requests without requiring full model retraining. By strategically partitioning data into distinct “shards” linked to different model versions, the SISA guideline allows for the targeted unlearning of specific data shards without requiring the participation of all shards. This method has been successfully applied in various domains, including unlearning in graph neural networks [chen2022graph] and large language models [chen2023unlearn]. However, SISA unlearning faces significant challenges in multi-client SL scenarios due to the interdependence between client model segments and the shared server model segment, violating the Isolated requirement of SISA. Consequently, the direct application of conventional SISA is impractical for existing SL scenarios.
@@
Label privacy under SL unlearning. Beyond unlearning, privacy concerns regarding intermediate outputs and labels in SL have been widely studied. U-shaped SL [GUPTA20181] avoids label propagation to the server by placing the output layer on the client side but violates SISA’s Isolated requirement by necessitating the return of intermediate values across all clients, disabling the independent unlearning and adding overhead to resource-limited clients. If labels are shared but privacy is required, existing research primarily addresses privacy risks by directly adding noise and obfuscating labels [ghazi2021deep, gao2024label] or further protecting intermediate outputs transmitted to the server. Wu et al. [wu2023federated] introduced a DP mechanism to perturb these transmitted outputs, while Li et al. [li2021label] developed the Marvell method to quantify privacy leakage and optimize noise structures. However, these methods overlook the exposure of sensitive information through label quantity and semantics. Xiao et al. [xiao2021mixing] obfuscated both labels and outputs via multiple linear combinations to reduce association with original data but still failed to conceal label quantity, potentially exposing sensitive information. Protecting the quantity of labels is critical in label-sharing SL, as revealing the exact number of labels allows adversaries to exploit supervised learning techniques that achieve higher accuracy by relying on this information, significantly increasing the risk of inference attacks and unauthorized model reconstruction.
@@
RRWithPrior [ghazi2021deep]
LPSC [gao2024label]
DP-based Defense [wu2023federated]
Marvell [li2021label]
MALM [xiao2021mixing]
U-shaped SL [GUPTA20181]
@@ 3.1 Split Learning
SL is a distributed ML technique designed to train neural networks without sharing raw data between $K$ clients, each owning datasets $D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$, and the server that holds no data [GUPTA20181, singh2019detailed, wu2023split]. A typical SL solution involves two key factors:
@@
The above processes enable training without sharing raw data or model details between clients and the server. In multi-client, single-server scenarios, two primary training methods are used: synchronizing model weights between clients after each epoch or alternating training epochs between clients and the server without synchronization [GUPTA20181].
Incremental training process. This process has been utilized by many learning frameworks, including FL [tan2022federated, nguyen2021federated] and SL, where a model is continuously updated by gradually incorporating new data over time. Each iteration of back-propagation during model training needs to be sent back to every client involved in the training process and builds incrementally on the previous ones, embodying the incremental nature of SL [9562559]. This process enables the model to dynamically learn and adapt as new data becomes available.
The challenges of applying SISA in multi-client SL scenarios stem from the incremental training process inherent in these setups. This iterative process undermines the isolated requirement of SISA due to the tight coupling between the server and client model segments through regular bidirectional propagation and the shared server model being incrementally trained across all clients. The knowledge to be unlearned, reflected in $W_{k}^{(t)}$, would have also influenced $W_{k^{\prime}}^{(t^{\prime})}$ for all $k^{\prime}\neq k$ and $t^{\prime}\geq t$, necessitating the recalculation of corresponding changes in subsequent updates, thereby making it impossible to apply SISA unlearning [liu2024decentralized, liu2024fishers].
@@ 3.2 Machine Unlearning
[…]
Retraining. The most direct and effective, though not efficient, method of implementing unlearning is retraining [retrain, retrain1]. In this method, the revoked sample is deleted, and the model is retrained on the original dataset minus the deleted sample, denoted as $D_{r}=D\setminus{((X_{i},Y_{q})}\subset D_{u})$. While this method is effective and straightforward, the computational overhead becomes prohibitive for complex models and large datasets.
Alternatives. Referring to fine-tuning, gradient ascent, and knowledge distillation, These unlearning methods are proposed to improve unlearning efficiency compared to retraining, but they still consider retraining as the baseline for unlearning effectiveness [qu2023learn, nguyen2022survey, distil1Unl], including:
@@
SISA. SISA [sisa] is a guideline designed to enable efficient and independent unlearning for those who request it, without affecting other participants. In SISA, the original training set $D_{o}$ is divided into $K$ disjoint shards $D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$}, each used to train a separate model $\mathcal{F}_{o}^{1},\mathcal{F}_{o}^{2},\ldots,\mathcal{F}_{o}^{K}$}. These models generate individual predictions, which are then aggregated to produce a global prediction. When an unlearning request is made, only the shard containing the specific data sample needs to be retrained, allowing the unlearning process to be conducted independently for the requesting party, without impacting other clients or the overall system.
@@ 4.1 Problem Definition
    G1: Independent unlearning. This goal ensures that unlearning affects only the requesting client and server, leaving other clients ($k’\neq k$) and their resources unaffected.
@@ 5 SplitWiper+: Privacy-Preserving Variation via Label Expansion
    Internal users. This includes clients ($\mathbb{C}_{inter}$) and servers ($\mathbb{S}_{inter}$) involved in training, modeled as honest-but-curious [li2021label]. They adhere to the protocol without tampering, such as sending incorrect features or causing failures. Each client ${C}_{inter}\in\mathbb{C}_{inter}$ knows the shared labels among participants but keeps their complete label set $\mathbb{Y}_{inter}$ private. Clients also aim to hide the quantity and semantics of their labels from the server $\mathbb{S}_{inter}$. This model emphasizes privacy preservation in internal label sharing.
@@
    External attackers. An external attacker $\mathbb{A}_{ext}$ is considered to know the distribution of the original dataset fed into the input layer at the targeted client $\mathbb{C}_{target}$, and the posterior distribution of the model at the server side. The attacker $\mathbb{A}_{ext}$ aims to use a shadow dataset that mimics these distributions to construct an attack model for conducting inference attacks [10.1145/3460120.3484756]. This player model focuses on the level of privacy preservation in the unlearning capability of the framework.
@@
Protecting label quantity matters. In G3, protecting the label privacy goes beyond protecting their semantics to include the quantity of labels, which is crucial for data privacy. By concealing the exact number of labels, adversaries are prevented from applying supervised learning methods that rely on knowing the number of classes for high accuracy. Instead, they are forced to use unsupervised learning techniques, lacking clear objectives and explicit class labels and generally resulting in lower accuracy [osisanwo2017supervised, love2002comparing]. Protecting label quantity thus reduces the risk of unauthorized model reconstruction and inference attacks, strengthening the overall privacy and security of the learning process.
@@ 5.1 Label Expansion in SplitWiper+
Label expansion and anonymization (Line 13-16). Once the complete set $\mathbb{Y}$ is aggregated, clients collaborate to assign an expansion factor $\gamma_{q}$ to each label $Y_{q}$. This factor, $\gamma_{q}$, can be a randomly selected integer within a predefined range $\Gamma_{p}$ for each $Y_{q}$, or be a constant integer $\gamma>1$ and uniformly applied across all labels, i.e., $\gamma_{q}=\gamma,\forall q$. Each label, $Y_{q}$, is then expanded to $\gamma_{q}$ instances, forming the initial expanded label set $\mathbb{Y}_{exp}=\{Y_{1}^{*},\cdots,Y_{e}^{*},\cdots Y_{E}^{*}\}$, where $E=\sum_{q=1}^{Q}{\gamma_{q}}$. This expansion process masks the total count $Q$ of real labels, preventing the server from inferring specific client tasks or data types based on label quantity. The elements within $\mathbb{Y}_{exp}$ are shuffled and re-indexed to anonymize label semantics. While we use a basic pseudonymization technique, our framework can also accommodate methods like hashing [ribeiro2019privacy] or encryption [zhang2020batchcrypt, deshpande2021sypse] based on specific needs. The mapping between $\mathbb{Y}$ and $\mathbb{Y}_{exp}$ is preserved in $\mathcal{G}_{Y}$, ensuring that labels are expanded and anonymized without losing semantic integrity.